OpenStudy (anonymous):

[4_01]Which of the following is the slope between the points (-6, 0) and (2, -9)? - -

OpenStudy (anonymous):

slope formula is \[\frac{ y2-y1 }{ x2-x1 }\] 2 is y2 0 is y1 2 is x2 -6 is x1

OpenStudy (anonymous):

\[\frac{ 2-0 }{ 2-(-6) }\]

OpenStudy (anonymous):

whered u get a 0?

OpenStudy (anonymous):

i meant another 2

OpenStudy (anonymous):

oops my bad

OpenStudy (anonymous):

do u have a kik or soemthing?

OpenStudy (anonymous):

it would look like this\[\frac{ -9-0 }{ 2-(-6) }\]

OpenStudy (anonymous):

-9/8

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

can u help me with more?

OpenStudy (anonymous):

ill see what I can do :)

OpenStudy (anonymous):

do you have kik ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

4_06]Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

OpenStudy (anonymous):

im not suree on this one

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

4_05]Choose the equation of the vertical line passing through the point (-4, 2). y = -4 x = -4 y = 2 x = 2 --------------------------------------------------------------------------------

OpenStudy (anonymous):

asnweer?

OpenStudy (anonymous):

Why are you throwing out more questions before the earlier one was answered? Answers are not given out here. We can help you learn how to solve it on your own. Read the code of conduct: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

4.06: Use the point-slope formula for a line: y-y1=m(x-x1) m is the slope of your line. (x1, y1) is any point on your line.

OpenStudy (anonymous):

4.05 Draw a vertical line through (-4,2). Study that and you should figure out what the equation of that line is. Either x or y is always equal to the same number at any point on your line.